• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 16:04:17 on January13,2002

    Entry number 80269

    keyword=day shift summary

    08:00 shift crew: J. Calarco, M. Roedelbronn, R. Michaels
    ----------------------------------------

    Run Summary
    ============

    Run...............Events...............Comments
    1478...............12.5 K..............15 cm Rad, LH2, OTR in
    1479................68 K.................same but OTR out
    1480...............163 K................same as 1479
    1481................28 K.................15 cm LH2, no radiator. (note: the halog start
    ..................................................run entry is wrong for run 1481.. no R).
    1482...............29 K..................same.
    1483...............45 K..................same but increased to 5 micoamps
    1484...............28 K..................same but now RCS magnet to zero. (Previously
    ..................................................RCS magnet had been 500 Amps).
    1485...............24 K...................same as 1484.
    1486...............30 K...................RCS magnet = 0 (zero) Amps. Note: RCS calor.
    ..................................................HV was changed at start of run 1486. Will repeat.
    1487...............30 K...................Good run with RCS magnet at 250 Amps.
    ..................................................P(hrs) = 1.711 and Angle(hrs) = 33.97
    1488...............38 K...................Good run with RCS magnet at 0 (zero) Amps.
    ..................................................P(hrs) = 1.711 and Angle (hrs) = 33.97
    Next we change momentum and angle. We cycle quads 2 and 3.
    1489...............222 K...................P(hrs) = 1.768 and Angle(hrs) = 34.59. RCS magnet
    ..................................................at zero Amps. 15 cm LH2 with no radiator.
    Another field and angle change:
    1490...............427 K...................P(hrs) = 1.654 and Angle(hrs) = 33.40, magnet zero.
    Another field and angle change:
    1491..............486 K...................P(hrs) = 1.556 and Angle = 35.68, magnet zero.
    1492.......Turn on RCS magnet to 250 Amps.. otherwise, same as 1492
    ...............Leave this for next shift...

    Comments
    =========

    1. We see lots of background, probably because beam is far off its
    proper path to accomodate the shifted target (see owl shift halogs).

    2. Some adjustments of HV by Gene, approx 10:30 to 12:30.

    3. 12:30 - Fixed some problems with datamon's deadtime calculation.
    A problem remains that it doesn't read T7, but if ps7=big, this is a small error.

    4. Alan Nathan says that the RCS magnet polarity is backwards.

    5. Target problems are severe. We cannot move it and so on.